8YBXMD7 7. You drill 375 frames but only 371 are defect-free. What is your percent defective? 1.33% 98.9% 1.07% O 1.1% 8. You 7&8
Answer #1

Answer

(7) Total number of frames = 375 and defect frames = 371

Number of defective frames = total frames - defect free frames = 375 - 371 = 4

we know that percent = [(number of defective frames)/(total)]*100

putting the values

Required percent = (4/375)*100 = 1.07%

Option C

(8) we know that gross pay = (hourly rate*number of hours at base wages)+(hourly rate*number of hours at prevailing wages)

From question, it is clear that

hourly rate for base wages = $10.25

number of hours at base wage = 32 hours

hourly rate for prevailing wages = $40

number of hours at prevailing wages = 8 hours

putting the values

we get

gross pay = ($10.25*32)+($40*8) = $328 + $320

So, gross pay = $648

option C